# Hemlet: A Heterogeneous Compute-in-Memory Chiplet Architecture for Vision Transformers with Group-Level Parallelism thanks: *Equal contribution

### v1

@@ IV-A Group-Level Parallelism: Motivation
IV-B Cross-layer Interleaved Distribution
@@
$\mathcal{C}_{j}=\{\mathbf{W}^{0}[:,j],\,\mathbf{W}^{1}[:,j],\,\dots,\,\mathbf{W}^{L-1}[:,j]\}.$ (1)
@@
Figure 6: Illustration of the interleaved weight distribution strategy. The laye-wise mapping method (left) statically assigns each group to a specific layer, while the interleaved approach (right) distributes weights from multiple layers across groups to enable concurrent group activation.
@@
$\mathrm{Interleave}\!\bigl(\mathbf{W}^{(1)},\,\mathbf{W}^{(2)},\dots,\mathbf{W}^{(x)}\bigr)\;=\;\mathbf{W}^{\mathrm{aug}}\in\mathbb{R}^{C_{\text{in}}\times xC_{\text{out}}}.$ (2)
Figure 7: Construction of the augmented weight matrix for each GLP_LayerSet
IV-C Deployment Approach
@@
$\mathbf{W}_{1}$ $\rightarrow\{\mathbf{W}1_{0},\mathbf{W}1_{1},\dots,\mathbf{W}1_{(k-1)}\},\quad\mathbf{W}1_{i}\in\mathbb{R}^{d\times d}$ (3)
@@
V System Dataflow Optimization
As introduced in Section III, Hemlet leverages heterogeneous CIM chiplets—ACIM for static VMMs and DCIM for runtime-generated dynamic VMMs. While this heterogeneous organization offers architectural flexibility, it also amplifies inter-chiplet data movement.
Figure 8: Dataflow Optimization. (a) Native dataflow for a Transformer block (basic layer-by-layer execution on heterogeneous chiplets); (b) Communication/computation pipeline (utilize the hierarchical buffers to overlap communication latency); (c) Blocked dynamic VMMs on DCIM chiplets (the standard Softmax is split into two stages so each intermediate $\mathbf{P}^{\prime}$ block can be consumed immediately, followed by a final normalization across blocks); (d) Optimized system dataflow (reduced inter-chiplet communication and improved utilization).
@@ V System Dataflow Optimization
For the dynamic VMMs operation, the input matrices $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are partitioned into smaller blocks along the sequence length for tiled processing considering the DCIM capacity limit, adopting the same block size $B_{L}$ as used in the previous stage to match the processing granularity. For the in-memory VMMs, blocks of $\mathbf{Q}$ and $\mathbf{V}$ will be written into the DCIM array as weights for $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ operations separately. To reduce the forward and backward transfers of the intermediate matrix $\mathbf{P}^{\prime}$ between the DCIM chiplet and the IDP, we adopt a placement strategy that prioritizes mapping the PEs responsible for both $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ of the same attention head to the same DCIM chiplet. While the process is tiled, fully executing the Softmax operation between $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ on local DCIM chiplets requires a much larger chiplet buffer to store the entire intermediate matrix $\mathbf{P}^{\prime}$ of size $B_{L}\times L$. To reduce this buffer overhead, we adopt an optimization inspired by FlashAttention [flashattention], a local $\mathrm{Softmax}_{\text{local}}$ is applied to each block result of $\mathbf{P^{\prime}}=\mathbf{Q}\mathbf{K}^{T}$, yielding $\mathbf{P^{\prime\prime}}=\mathrm{Softmax}_{\text{local}}(\mathbf{P^{\prime}})$. The corresponding block output is then obtained as 2 $\mathbf{S^{\prime}}=\mathbf{P^{\prime\prime}}\mathbf{V}$ in Fig. 8 (c). After collecting all $\mathbf{S^{\prime}}$ blocks along the standard Softmax dimension, a 3 global normalization is performed to produce the $\mathbf{S}$ block. In this way, each $\mathbf{P}^{\prime}$ block can be consumed immediately, avoiding much larger on-chip buffer of the DCIM chiplet or the forward and backward transfers between DCIM chiplets and IDP chiplet.
VI Evaluation
VI-A Experimental Methodology
@@
Simulator and Modeling Framework: We develop a custom cycle-accurate simulator tailored for heterogeneous CIM-based chiplet systems. The simulation flow integrates hardware-level modeling, dataflow-level scheduling, and event-driven instruction execution across computations on chiplets and inter-chiplet communication. For ACIM chiplets, we adopt NeuroSim [NeuroSimv1.3] to model RRAM-based subarrays and other digital blocks at a 22 nm technology node. DCIM chiplets and the IDP are also extended from NeuroSim and assumed at the 7 nm technology node, with DCIM array performance from the AutoDCIM design [autodcim] and scaled to 7 nm using DeepScale [sarangi2021deepscaletool]. To estimate inter-chiplet communication cost, we extend BookSim [jiang2013detailed] for latency estimation under a mesh topology. The corresponding energy consumption is evaluated using post-silicon data from Simba [shao2019simba]. To determine the hardware configuration of ACIM, we utilize MICSim [micsim] to evaluate ImageNet-1K[imagenet] classification performance under the ViT-B/16 model. We perform quantization-aware training (QAT) with I-ViT[li2023ivit], achieving $80\%$ accuracy. Our results indicate that with an ADC precision of 9 bits and 2-bit RRAM with an $R{on}/R_{off}$ ratio of 150[rram150], no accuracy degradation is observed compared to our software baseline.
@@
System Configuration: The system is configured to ensure all pretrained weights are fully mapped into ACIM chiplets before inference, consistent with prior works [retransformer, x-former, shafiee2016isaac, chi2016prime, siam]. An ACIM chiplet consists of multiple tiles. Each tile features a fixed configuration of 16 PEs, with each PE comprising 6 ACIM subarrays. The group size for the ACIM subarray is set to 8, denoted as G8. Meanwhile, the DCIM chiplet comprises multiple PEs, with each PE integrating 4 DCIM subarrays. Unless otherwise specified, we default to an ACIM chiplet with 16 Tiles and a DCIM chiplet with 16 PEs; this default configuration is denoted as T16P16. The default NoP bandwidth is set to 256 GB/s per chiplet. To account for the varying capacities of chiplets, we assume that the number of tiles for the ACIM chiplet and the number of PEs for the DCIM chiplet may vary. To analyze the impact of chiplet size and communication bandwidth, we evaluate configurations of T9P9, T16P16, and T25P25 under a range of NoP bandwidths from 64 GB/s to 256 GB/s. The NoP bandwidth per chiplet represents the total bandwidth of all inter-chiplet links, specifically the combined bandwidth of eight unidirectional links configured in a mesh topology.
TABLE I: Vision Transformer Model Configuration
@@
Figure 9: The GLP speedup breakdown of ACIM chiplet
Workloads: We evaluate our system using standard ViT models [vit], covering three representative scales: ViT-S/16, ViT-B/16, and ViT-L/16. These models span a wide range of parameter counts and computational requirements, allowing us to assess the scalability and efficiency of our proposed architecture and GLP strategy. All models are evaluated under the image resolution of $224\times 224$, using $16\times 16$ patching, and are fully quantized to 8-bit with I-ViT [li2023ivit] on the ImageNet-1K[imagenet]. The detailed model parameters, including embedding dimension, number of Transformer blocks, and the Sequence length, are summarized in Table I.
VI-B Impact of GLP on ACIM Chiplet Execution
@@
Overall, GLP achieves consistent performance improvements across all settings, primarily driven by improvements in subarray execution. This outcome aligns with the design objective of GLP, which interleaves weights across memory groups to improve the SA level parallelism. In addition, both buffer and IC exhibit performance improvements under GLP mapping. This benefit stems from the weight column-wise interleaving strategy in GLP, which activates more groups and thereby enables concurrent operation of subarrays, PEs, tiles, and even chiplets. As a result, the output data are more evenly distributed across each hierarchy, which in turn reduces buffer pressure and shortens communication latency.
@@
Figure 10: System performance analysis under different configurations
VI-C System-level Performance Analysis
@@
Ablation Study: To evaluate the effects of the proposed GLP mapping and system dataflow optimization on the system performance, we perform an ablation study across different hardware settings. The analysis considers three execution strategies: Baseline, which adopts the conventional layer-wise mapping with the native system dataflow described in Section V; Baseline + GLP, which further applies the proposed GLP mapping; and Hemlet, which incorporates both GLP and system-level dataflow optimization.
@@
Fig. 10 (a) shows that GLP reduces latency in most settings by improving ACIM throughput, with the maximum system speedup reaching $2.55\times$. However, under small chiplet sizes (T9P9) and very low bandwidth (64 GB/s), it can even fall below the baseline, as NoP communication becomes the dominant bottleneck. The degradation occurs because GLP tends to distribute the weights of a single layer into more groups. Consequently, compared to conventional methods, GLP often allocates more chiplets for a single layer when chiplet capacity is low. This results in an increased number of active sources and destinations on the NoP during inference, leading to higher traffic volume and greater communication complexity. Under conditions of limited bandwidth, this can cause significant NoP congestion, resulting in throughput degradation. However, when bandwidth is sufficient, these issues are mitigated. The system-level dataflow optimization further improves performance by alleviating inter-chiplet communication, providing greater benefits especially under conditions of limited bandwidth and smaller chiplet sizes. Overall, the proposed Hemlet architecture, which integrates GLP and system-level dataflow optimization, ensures consistent acceleration across all tested configurations, delivering system-level speedups ranging from $1.44\times$ to $4.07\times$.
@@
As shown in Fig.10 (b), we can observe that GLP mapping introduces extra energy overhead in all settings. This arises because, under GLP, the weights of each layer are interleaved across more groups, leaving each ACIM subarray with fewer weight columns. As a result, each input is reused less within one subarray and must be loaded into more subarrays, thereby increasing data movement and overall energy consumption. Nevertheless, the additional energy overhead remains within about $15\%$. After applying the system-level dataflow optimization, the overall energy consumption can be reduced. This is because the optimization decreases inter-chiplet communication, thereby lowering the energy cost of data transfers. As a result, the system-level dataflow optimization can partially compensate for the additional energy overhead introduced by GLP.
@@
Comparison with Existing Work: In Table II, we compare our design with two state-of-the-art ViT accelerators. The work presented in [dac24dcim] utilizes DCIM arrays as the computation cores for both static VMMs and dynamic VMMs. While the reload capability of the DCIM computation units allows for a monolithic design with limited on-chip resources, this approach results in relatively low throughput due to the necessary weight reloads. On the other hand, H3DAtten [li2023h3datten] shares the same design principles as Hemlet, utilizing heterogeneous computing units and employing ACIM arrays to store all weights on-chip. To address area constraints, they employ 3D integration of the eNVM arrays with the other digital circuits. This 3D integration mitigates technology lock-in, as the top and bottom tiers can be fabricated using different process nodes. Hemlet achieves superior throughput compared to H3DAtten, which can be mainly attributed to two key reasons. First, H3DAtten employs a lower precision ADC, necessitating a reduction in row parallelism in the CIM computation to maintain accuracy. Second, Hemlet enhances throughput further by implementing GLP and optimizing dataflow at the system-level. H3DAtten demonstrates better energy efficiency than Hemlet, which can also be attributed to two primary factors. First, the 3D integration in H3DAtten allows for a simpler NoC and incurs lower communication overhead compared to the NoP adopted by Hemlet. Second, H3DAtten locates its ADCs on the bottom tier and fabricates them using a more advanced technology node. In contrast, Hemlet integrates the ADCs on the ACIM chiplet, sharing the same technology node as the eNVM arrays. While this 3D integration in H3DAtten enhances energy efficiency, it also introduces greater design complexity and reduces flexibility due to the tight coupling between the top-tier array and the bottom-tier ADC.
TABLE II: Comparison of recent CIM accelerator designs.
Work DAC’24[dac24dcim] H3DAtten[li2023h3datten] Hemlet
@@
Throughput 0.828 1.61 8.68
Energy Efficiency N/A 7.1 3.86

### v2

@@ 4.1. Group-Level Parallelism: Motivation
4.2. Cross-layer Interleaved Distribution
@@
(1) $\mathcal{C}_{j}=\{\mathbf{W}^{0}[:,j],\,\mathbf{W}^{1}[:,j],\,\dots,\,\mathbf{W}^{L-1}[:,j]\}.$
@@
Figure 6. Illustration of the interleaved weight distribution strategy. The laye-wise mapping method (left) statically assigns each group to a specific layer, while the interleaved approach (right) distributes weights from multiple layers across groups to enable concurrent group activation.
@@
(2) $\mathrm{Interleave}\!\bigl(\mathbf{W}^{(1)},\,\mathbf{W}^{(2)},\dots,\mathbf{W}^{(x)}\bigr)\;=\;\mathbf{W}^{\mathrm{aug}}\in\mathbb{R}^{C_{\text{in}}\times xC_{\text{out}}}.$
Figure 7. Construction of the augmented weight matrix for each GLP_LayerSet
4.3. Deployment Approach
@@
(3) $\mathbf{W}_{1}$ $\rightarrow\{\mathbf{W}1_{0},\mathbf{W}1_{1},\dots,\mathbf{W}1_{(k-1)}\},\quad\mathbf{W}1_{i}\in\mathbb{R}^{d\times d}$
@@
5. System Dataflow Optimization
As introduced in Section 3, Hemlet leverages heterogeneous CIM chiplets—ACIM for static VMMs and DCIM for runtime-generated dynamic VMMs. While this heterogeneous organization offers architectural flexibility, it also amplifies inter-chiplet data movement.
Figure 8. Dataflow Optimization. (a) Native dataflow for a Transformer block (basic layer-by-layer execution on heterogeneous chiplets); (b) Communication/computation pipeline (utilize the hierarchical buffers to overlap communication latency); (c) Blocked dynamic VMMs on DCIM chiplets (the standard Softmax is split into two stages so each intermediate $\mathbf{P}^{\prime}$ block can be consumed immediately, followed by a final normalization across blocks); (d) Optimized system dataflow (reduced inter-chiplet communication and improved utilization).
@@ 5. System Dataflow Optimization
For the dynamic VMMs operation, the input matrices $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are partitioned into smaller blocks along the sequence length for blocked processing considering the DCIM capacity limit, adopting the same block size $B_{L}$ as used in the previous stage to match the processing granularity. For the in-memory VMMs, blocks of $\mathbf{Q}$ and $\mathbf{V}$ will be written into the DCIM array as weights for $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ operations separately. To reduce the forward and backward transfers of the intermediate matrix $\mathbf{P}^{\prime}$ between the DCIM chiplet and the IDP, we adopt a placement strategy that prioritizes mapping the PEs responsible for both $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ of the same attention head to the same DCIM chiplet. While the process is blocked, fully executing the Softmax operation between $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ on local DCIM chiplets requires a much larger chiplet buffer to store the entire intermediate matrix $\mathbf{P}^{\prime}$ of size $B_{L}\times L$. To reduce this buffer overhead, we adopt an optimization inspired by FlashAttention (flashattention), a local $\mathrm{Softmax}_{\text{local}}$ is applied to each block result of $\mathbf{P^{\prime}}=\mathbf{Q}\mathbf{K}^{T}$, yielding $\mathbf{P^{\prime\prime}}=\mathrm{Softmax}_{\text{local}}(\mathbf{P^{\prime}})$. The corresponding block output is then obtained as 2 $\mathbf{S^{\prime}}=\mathbf{P^{\prime\prime}}\mathbf{V}$ in Fig. 8 (c). After collecting all $\mathbf{S^{\prime}}$ blocks along the standard Softmax dimension, a 3 global normalization is performed to produce the $\mathbf{S}$ block. In this way, each $\mathbf{P}^{\prime}$ block can be consumed immediately, avoiding much larger on-chip buffer of the DCIM chiplet or the forward and backward transfers between DCIM chiplets and IDP chiplet.
6. Evaluation
6.1. Experimental Methodology
@@
Simulator and Modeling Framework: We develop a custom cycle-accurate simulator tailored for heterogeneous CIM-based chiplet systems. The simulation flow integrates hardware-level modeling, dataflow-level scheduling, and event-driven instruction execution across computations on chiplets and inter-chiplet communication. For ACIM chiplets, we adopt NeuroSim (NeuroSimv1.3) to model RRAM-based subarrays and other digital blocks at a 22 nm technology node. DCIM chiplets and the IDP are also extended from NeuroSim and assumed at the 7 nm technology node, with DCIM array performance from the AutoDCIM design (autodcim) and scaled to 7 nm using DeepScale (sarangi2021deepscaletool). To estimate inter-chiplet communication cost, we extend BookSim (jiang2013detailed) for latency estimation under a mesh topology. The corresponding energy consumption is evaluated using post-silicon data from Simba (shao2019simba). To determine the hardware configuration of ACIM, we utilize MICSim (micsim) to evaluate ImageNet-1K(imagenet) classification performance under the ViT-B/16 model. We perform quantization-aware training (QAT) with I-ViT(li2023ivit), achieving $80\%$ accuracy. Our results indicate that with an ADC precision of 9 bits and 2-bit RRAM with an $R{on}/R_{off}$ ratio of 150(rram150), no accuracy degradation is observed compared to our software baseline.
@@
System Configuration: The system is configured to ensure all pretrained weights are fully mapped into ACIM chiplets before inference, consistent with prior works (retransformer; x-former; shafiee2016isaac; chi2016prime; siam). An ACIM chiplet consists of multiple ACIM PEs. Each ACIM PE features a fixed configuration of $60$ ACIM subarrays ($128\times 128$). The group size for the ACIM subarray is set to 8. Meanwhile, the DCIM chiplet comprises multiple DCIM PEs, with each DCIM PE integrating 4 DCIM subarrays ($64\times 64$). To account for the varying capacities of chiplets, we assume that the number of ACIM PEs for the ACIM chiplet and the number of DCIM PEs for the DCIM chiplet may vary. To analyze the impact of chiplet size and communication bandwidth, we evaluate configurations of A18D9 (each ACIM chiplet with 18 ACIM PEs and each DCIM chiplet with 9 DCIM PEs), A32D16, and A50D25 under a range of NoP bandwidths from 8 GB/s to 32 GB/s. The NoP communication network is configured in a 2D mesh topology, which follows the design in Simba(shao2019simba).
Table 1. Vision Transformer Model Configuration
@@
Figure 9. The GLP speedup breakdown of ACIM chiplet
Workloads: We evaluate our system using standard ViT models (vit), covering three representative scales: ViT-S/16, ViT-B/16, and ViT-L/16. These models span a wide range of parameter counts and computational requirements, allowing us to assess the scalability and efficiency of our proposed architecture and GLP strategy. All models are evaluated under the image resolution of $224\times 224$, using $16\times 16$ patching, and are fully quantized to 8-bit with I-ViT (li2023ivit) on the ImageNet-1K(imagenet). The detailed model parameters, including embedding dimension, number of Transformer blocks, and the Sequence length, are summarized in Table 1.
6.2. Impact of GLP on ACIM Chiplet Execution
@@
Overall, GLP achieves consistent performance improvements across all settings, primarily driven by improvements in subarray execution. This outcome aligns with the design objective of GLP, which interleaves weights across memory groups to improve the SA level parallelism. In addition, both buffer and IC exhibit performance improvements under GLP mapping. This benefit stems from the weight column-wise interleaving strategy in GLP, which activates more groups and thereby enables concurrent operation of subarrays, PEs, and even chiplets. As a result, the output data are more evenly distributed across each hierarchy, which in turn reduces buffer pressure and shortens communication latency.
@@
Figure 10. Normalized system latency and energy under different chiplet configurations and NoP bandwidths (8/16/32 GB/s).
6.3. System-level Performance Analysis
@@
Ablation Study: To evaluate the effects of the proposed GLP mapping and system dataflow optimization on the system performance, we perform an ablation study across different hardware settings. The analysis considers three execution strategies: Baseline, which adopts the conventional layer-wise mapping with the native system dataflow described in Section 5; Baseline + GLP, which further applies the proposed GLP mapping; and Hemlet, which incorporates both GLP and system-level dataflow optimization.
@@
Fig. 10(a) shows that GLP reduces latency in most settings by improving ACIM throughput, with the maximum system speedup reaching $2.53\times$. The effectiveness of this improvement is senstive on both NoP bandwidth and chiplet size. Under lower bandwidth NoP configurations, inter-chiplet communication occupies a larger fraction of the overall execution time, which limits the achievable benefit from ACIM-side acceleration. As the NoP bandwidth increases, the communication bottleneck is gradually alleviated, allowing the throughput advantages of GLP-enabled ACIM execution to translate more effectively into system speedup. The system-level dataflow optimization further improves performance by alleviating inter-chiplet communication, providing greater benefits especially under conditions of limited bandwidth and smaller chiplet sizes. Overall, the proposed Hemlet architecture, which integrates GLP and system-level dataflow optimization, ensures consistent acceleration across all tested configurations, delivering system-level speedups ranging from $1.89\times$ to $4.47\times$.
@@
As shown in Fig.10 (b), we can observe that GLP mapping introduces extra energy overhead in all settings. This arises because, under GLP, the weights of each layer are interleaved across more groups, leaving each ACIM subarray with fewer weight columns. As a result, each input is reused less within one subarray and must be loaded into more subarrays, thereby increasing data movement and overall energy consumption. After applying the system-level dataflow optimization, the overall energy consumption can be reduced. This is because the optimization decreases inter-chiplet communication, thereby lowering the energy cost of data transfers. As a result, the system-level dataflow optimization can partially compensate for the additional energy overhead introduced by GLP. Still, the remaining energy overhead is around $10\%$.
@@
Comparison with Existing Work: In Table 2, we compare our design with two state-of-the-art ViT accelerators. The reported results for Hemlet correspond to the configuration using the A32D16 chiplet setup with 32 GB/s NoP bandwidth, evaluated on the ViT-L/16 model for end-to-end inference, where each processed element in non-VMM operations is counted as one operation. The work presented in (dac24dcim) utilizes DCIM arrays as the computation cores for both static VMMs and dynamic VMMs. While the reload capability of the DCIM computation units allows for a monolithic design with limited on-chip resources, this approach results in relatively low throughput due to the necessary weight reloads. On the other hand, H3DAtten (li2023h3datten) shares the same design principles as Hemlet, utilizing heterogeneous computing units and employing ACIM arrays to store all weights on-chip. To address area constraints, they employ 3D integration of the eNVM arrays with the other digital circuits. This 3D integration mitigates technology lock-in, as the top and bottom tiers can be fabricated using different process nodes. Hemlet achieves superior throughput compared to H3DAtten, which can be mainly attributed to two key reasons. First, H3DAtten employs a lower precision ADC, necessitating a reduction in row parallelism in the CIM computation to maintain accuracy. Second, Hemlet enhances throughput further by implementing GLP and optimizing dataflow at the system-level. H3DAtten demonstrates better energy efficiency than Hemlet, which can also be attributed to two primary factors. First, the 3D integration in H3DAtten allows for a simpler NoC and incurs lower communication overhead compared to the NoP adopted by Hemlet. Second, H3DAtten locates its ADCs on the bottom tier and fabricates them using a more advanced technology node. In contrast, Hemlet integrates the ADCs on the ACIM chiplet, sharing the same technology node as the eNVM arrays. While this 3D integration in H3DAtten enhances energy efficiency, it also introduces greater design complexity and reduces flexibility due to the tight coupling between the top-tier array and the bottom-tier ADC.
Table 2. Comparison of recent CIM accelerator designs.
Work DAC’24(dac24dcim) H3DAtten(li2023h3datten) Hemlet
@@
Throughput 0.828 1.61 9.24
Energy Efficiency N/A 7.1 4.98